Reports 1-1 of 1 Clear search Modify search
AdV-ALS (Green sources)
derossi, gosselin, valentini - 17:49 Thursday 10 June 2021 (52112) Print this report
realignment of the green beams on the EIB

Since the green beam coming from the North Arm was quite misaligned, this afternoon we went in the laser lab to optimize the coupling into the fiber.

We blocked the EIB, Michele locked the arms with the green and we optimized by adjusting the tilt of the mirrors just before the collimators (see attached plot).

Images attached to this report
Search Help